A chemist places 0.550 g KCl in a flask and adds water until the total volume is 0.900 L. Calculate the molarity of the solution.

What is the concentration in millimolar (mM)?


1
Expert's answer
2022-03-30T02:07:01-0400

No of mole =mass/molar mass. Molar mass of KCl= 74.55g/mol

=0.550g/74.55

= 0.0074mole

Concentration = mole/volume = 0.0074/0.900

=8.2 × 10-3 M = 8.2 × 10-6 mM